# Khushbu Pahwa3, Jian Ding1, Mohamed Elhoseiny1

### v4

@@ B.3 Data Copyrights
Our benchmark provides only annotations and uses two widely adopted, publicly available video datasets that provide legally compliant, downsampled video content (e.g., 3 FPS, scene shots). Specifically, we use: (1) six full-season TV shows from the TVQA Lei et al. (2019) dataset, based on the long-form version in GoldFish Ataallah et al. (2024b), and (2) movies from the MovieNet Huang et al. (2020) dataset. For the TV shows, transcripts were sourced from reputable platforms such as foreverdreaming and manually verified by the authors for alignment with the video. This usage is consistent with fair use for academic research. For movies, we use transcripts provided by MovieNet Huang et al. (2020), which are already filtered and accurately aligned with dialogue and subtitles.
@@ B.5 Duplications Filtering
We first embedded each question and its answer choices using M3-Embedding (Chen et al., 2024a), then computed pairwise cosine similarity to detect potential duplicates. We experimented with similarity thresholds of 90%, 95%, and 98%. As shown in Figure 5, the vast majority of questions are unique across all skill categories. Only two skills—Temporal Questions and Character Actions—produced potential duplicates. However, upon manual inspection, these were found to be false positives. For example, the following two questions were flagged as duplicates due to high embedding similarity, despite a critical semantic difference:
@@ B.6 Human Verification
The collected timestamps were cross-referenced with the dataset to validate both the extracted events and their chronological consistency. This also enabled the generation of temporal certificates for each Q/A pair, following the methodology introduced in EgoSchema Mangalam et al. (2023). A temporal certificate represents the number of video segments required to answer a given question. Following Mangalam et al. (2023), video clips shorter than 0.1 seconds were discarded, and consecutive segments within a 5-second window were merged into a single clip. The total duration of these merged segments defines the certificate length.
@@ B.7 Models Configuration
GPT-4o OpenAI (2024): GPT-4o does not support direct ingestion of .mp4 video files. To accommodate this, we sample up to 450 frames per video and provide the model with the corresponding subtitle text and question.
Gemini-Flash 2.0 Gemini (2024): Developed by Google, Gemini-Flash 2.0 supports native video input. For TV shows, we supply the model with the full video and the associated question. For movies—which lack audio—we additionally provide the subtitle file alongside the video and question.
Qwen2.5-VL Bai et al. (2025) and Qwen2-VL Wang et al. (2024a): These models accept up to 768 input frames, fitting within the memory limits of an NVIDIA A100 GPU (80 GB) under the authors’ recommended inference settings. Subtitles are concatenated with the input question to provide multimodal context.
Video-Chat-Flash Li et al. (2025): Although the model supports up to 10,000 frames, we limit input to 1,000 frames due to memory constraints on the A100 GPU (80 GB). Other configurations follow the default setup. Subtitles are concatenated with the input question to enrich contextual understanding.
InternVL3.0 Zhu et al. (2025a), InternVL2.5 Chen et al. (2024b), and InternVL2 InternVL (2): For the InternVL family, we evaluate models using both 16 and 128 frames, the latter being the maximum supported without hallucination on an A100 GPU. Although these models can technically process more frames, exceeding 128 often introduces unstable outputs. Subtitles corresponding to the selected frames are appended to the input question.
LLaVA-OneVision Li et al. (2024a): This model supports both image and video inputs. We evaluate performance at 16 and 128 frames—found to be the upper bound for stable performance on an A100 GPU. Subtitles for selected frames are appended to the question.
LongVU Shen et al. (2024): LongVU is evaluated using 512 frames, staying within the memory and context window limits of the A100 GPU. Subtitles aligned with the selected frames are appended to the input question.
InternLM-XComposer-2.5 Zhang et al. (2024): We use the default configuration with a 16-frame window and increase the number of clips per video to 120. Subtitles aligned with these clips are appended to the input question. All evaluations are conducted on an A100 GPU.
Goldfish and MiniGPT-4-Video Ataallah et al. (2024b): We use the Mistral backbone with default parameters, retrieving $k=3$ video clips per query. Each 80-second video yields 60 frames (approx. 1.3 FPS). Although the model supports up to 90 frames, we limit it to 60 to reserve space for the question, subtitles, and output tokens—reducing hallucinations due to context overflow. Subtitles are interleaved with the frames, following the MiniGPT-4-Video configuration. For MiniGPT-4-Video, we similarly sample 60 frames per video, interleave them with subtitles, and prompt the model with the question for generation.
@@ Appendix C Extended Related Work
Early work like Grounding Semantic Roles in Images (Silberer and Pinkal, 2018) and Grounded Situation Recognition (Pratt et al., 2020) showed how semantic roles can be aligned with visual regions in images. This concept was extended to videos by Video Object Grounding (Sadhu et al., 2020) and Visual Semantic Role Labeling (Sadhu et al., 2021), emphasizing semantic alignment for event understanding. Building on these ideas, HERO (Li et al., 2020) proposed a hierarchical encoder that fuses visual and textual modalities, setting the stage for video-language pretraining. Recent models like VSGR (Mao et al., 2022) further reinforce this trajectory by explicitly building video scene graphs to reason over semantic elements and their relations. In this context, our work introduces a benchmark that evaluates models’ ability to track and reason about evolving semantic roles across long-form, multimodal narratives.
@@
Recent advancements in commercial AI models have introduced native support for long-form video understanding. For instance, Google’s Gemini Flash family (Gemini, 2024) that provides a one million-token context window for coherent multi-hour understanding, and GPT-4o (OpenAI, 2024) that processes video frames in conjunction with subtitles within its 128K-token context window. In contrast to the commercial solutions, Open-source efforts rely on compression, memory, and retrieval techniques. Qwen2.5-VL (Bai et al., 2025) extends dynamic resolution temporally to spot key events in hours-long streams and ground them visually. Qwen2-VL (Wang et al., 2024a) uses Multimodal Rotary Position Embedding (M-RoPE) to encode up to 768 frames, preserving spatial and temporal structure. LongVU (Shen et al., 2024) applies cross-modal queries and inter-frame dependencies for adaptive compression, trimming redundancy while keeping salient content. InternLM-XComposer-2.5-OmniLive (Zhang et al., 2024) pairs a short-term memory buffer with on-the-fly compression into long-term storage for later retrieval. Goldfish (Ataallah et al., 2024b) segments videos into clips and retrieves the top-k most relevant, and LLaMA-VID (Li et al., 2023) represents each frame with only two tokens for maximal efficiency. VideoINSTA (Liao et al., 2024) enables long video understanding without pretraining by combining event-based temporal reasoning with LLMs and self-reflective spatial-temporal queries, achieving strong results on EgoSchema and NExT-QA. Despite these innovations, truly unrestricted long-form video understanding, with deep temporal dependencies and narrative complexity, remains an open challenge.
@@
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (Bai et al., 2025), Qwen2-VL (Wang et al., 2024a), and LongVU (Wu et al., 2024), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (Liang et al., 2024; Lei et al., 2021), action classification (Caba Heilbron et al., 2015; Carreira and Zisserman, 2017; Soomro et al., 2012), video reasoning (Xie et al., 2024; Xiao et al., 2021), and video QA (Xu et al., 2017; Lei et al., 2019; Jang et al., 2017). Among longer video benchmarks, TVQA (Lei et al., 2019) is an early example with over 152.5 K QA pairs across 21.8 K clips (1.27 min avg.), totaling 461.2 hours. More recent datasets like MovieChat-1K(Song et al., 2023) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(Ates et al., 2020) and GazeVQA (Ilaslan et al., 2023) extend to causal and embodied reasoning. Other long-form efforts include MLVU (Zhou et al., 2024), MoVQA (Zhang et al., 2023b), and Video-MME (Fu et al., 2024), though their average durations peak at 17 minutes. LVBench(Wang et al., 2024b) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(Mangalam et al., 2023) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count in terms of long video understanding (87.7K).

### v5

@@ B.3 Data Copyrights
Our benchmark provides only annotations and uses two widely adopted, publicly available video datasets that provide legally compliant, downsampled video content (e.g., 3 FPS, scene shots). Specifically, we use: (1) six full-season TV shows from the TVQA lei2019tvqa dataset, based on the long-form version in GoldFish ataallah2024goldfishvisionlanguageunderstandingarbitrarily, and (2) movies from the MovieNet huang2020movienet dataset. For the TV shows, transcripts were sourced from reputable platforms such as foreverdreaming and manually verified by the authors for alignment with the video. This usage is consistent with fair use for academic research. For movies, we use transcripts provided by MovieNet huang2020movienet, which are already filtered and accurately aligned with dialogue and subtitles.
@@ B.5 Duplications Filtering
We first embedded each question and its answer choices using M3-Embedding (bge-m3), then computed pairwise cosine similarity to detect potential duplicates. We experimented with similarity thresholds of 90%, 95%, and 98%. As shown in Figure 5, the vast majority of questions are unique across all skill categories. Only two skills—Temporal Questions and Character Actions—produced potential duplicates. However, upon manual inspection, these were found to be false positives. For example, the following two questions were flagged as duplicates due to high embedding similarity, despite a critical semantic difference:
@@ B.6 Human Verification
The collected timestamps were cross-referenced with the dataset to validate both the extracted events and their chronological consistency. This also enabled the generation of temporal certificates for each Q/A pair, following the methodology introduced in EgoSchema mangalam2023egoschema. A temporal certificate represents the number of video segments required to answer a given question. Following mangalam2023egoschema, video clips shorter than 0.1 seconds were discarded, and consecutive segments within a 5-second window were merged into a single clip. The total duration of these merged segments defines the certificate length.
@@ B.7 Models Configuration
GPT-4o GPT-4o: GPT-4o does not support direct ingestion of .mp4 video files. To accommodate this, we sample up to 450 frames per video and provide the model with the corresponding subtitle text and question.
Gemini-Flash 2.0 gemini_flash: Developed by Google, Gemini-Flash 2.0 supports native video input. For TV shows, we supply the model with the full video and the associated question. For movies—which lack audio—we additionally provide the subtitle file alongside the video and question.
Qwen2.5-VL bai2025qwen25vltechnicalreport and Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels: These models accept up to 768 input frames, fitting within the memory limits of an NVIDIA A100 GPU (80 GB) under the authors’ recommended inference settings. Subtitles are concatenated with the input question to provide multimodal context.
Video-Chat-Flash li2025videochatflashhierarchicalcompressionlongcontext: Although the model supports up to 10,000 frames, we limit input to 1,000 frames due to memory constraints on the A100 GPU (80 GB). Other configurations follow the default setup. Subtitles are concatenated with the input question to enrich contextual understanding.
InternVL3.0 zhu2025internvl3exploringadvancedtraining, InternVL2.5 InternVL2.5, and InternVL2 InternVL2: For the InternVL family, we evaluate models using both 16 and 128 frames, the latter being the maximum supported without hallucination on an A100 GPU. Although these models can technically process more frames, exceeding 128 often introduces unstable outputs. Subtitles corresponding to the selected frames are appended to the input question.
LLaVA-OneVision li2024llavaonevisioneasyvisualtask: This model supports both image and video inputs. We evaluate performance at 16 and 128 frames—found to be the upper bound for stable performance on an A100 GPU. Subtitles for selected frames are appended to the question.
LongVU shen2024longvuspatiotemporaladaptivecompression: LongVU is evaluated using 512 frames, staying within the memory and context window limits of the A100 GPU. Subtitles aligned with the selected frames are appended to the input question.
InternLM-XComposer-2.5 zhang2024internlmxcomposer25versatilelargevision: We use the default configuration with a 16-frame window and increase the number of clips per video to 120. Subtitles aligned with these clips are appended to the input question. All evaluations are conducted on an A100 GPU.
Goldfish and MiniGPT-4-Video ataallah2024goldfishvisionlanguageunderstandingarbitrarily: We use the Mistral backbone with default parameters, retrieving $k=3$ video clips per query. Each 80-second video yields 60 frames (approx. 1.3 FPS). Although the model supports up to 90 frames, we limit it to 60 to reserve space for the question, subtitles, and output tokens—reducing hallucinations due to context overflow. Subtitles are interleaved with the frames, following the MiniGPT-4-Video configuration. For MiniGPT-4-Video, we similarly sample 60 frames per video, interleave them with subtitles, and prompt the model with the question for generation.
@@ Appendix C Extended Related Work
Early work like Grounding Semantic Roles in Images (silberer2018grounding) and Grounded Situation Recognition (pratt2020grounded) showed how semantic roles can be aligned with visual regions in images. This concept was extended to videos by Video Object Grounding (sadhu2020video) and Visual Semantic Role Labeling (sadhu2021visual), emphasizing semantic alignment for event understanding. Building on these ideas, HERO (li2020hero) proposed a hierarchical encoder that fuses visual and textual modalities, setting the stage for video-language pretraining. Recent models like VSGR (mao2022dynamic) further reinforce this trajectory by explicitly building video scene graphs to reason over semantic elements and their relations. In this context, our work introduces a benchmark that evaluates models’ ability to track and reason about evolving semantic roles across long-form, multimodal narratives.
@@
Recent advancements in commercial AI models have introduced native support for long-form video understanding. For instance, Google’s Gemini Flash family (gemini_flash) that provides a one million-token context window for coherent multi-hour understanding, and GPT-4o (GPT-4o) that processes video frames in conjunction with subtitles within its 128K-token context window. In contrast to the commercial solutions, Open-source efforts rely on compression, memory, and retrieval techniques. Qwen2.5-VL (bai2025qwen25vltechnicalreport) extends dynamic resolution temporally to spot key events in hours-long streams and ground them visually. Qwen2-VL (wang2024qwen2vlenhancingvisionlanguagemodels) uses Multimodal Rotary Position Embedding (M-RoPE) to encode up to 768 frames, preserving spatial and temporal structure. LongVU (shen2024longvuspatiotemporaladaptivecompression) applies cross-modal queries and inter-frame dependencies for adaptive compression, trimming redundancy while keeping salient content. InternLM-XComposer-2.5-OmniLive (zhang2024internlmxcomposer25versatilelargevision) pairs a short-term memory buffer with on-the-fly compression into long-term storage for later retrieval. Goldfish (ataallah2024goldfishvisionlanguageunderstandingarbitrarily) segments videos into clips and retrieves the top-k most relevant, and LLaMA-VID (li2023llamavid) represents each frame with only two tokens for maximal efficiency. VideoINSTA (liao2024videoinsta) enables long video understanding without pretraining by combining event-based temporal reasoning with LLMs and self-reflective spatial-temporal queries, achieving strong results on EgoSchema and NExT-QA. Despite these innovations, truly unrestricted long-form video understanding, with deep temporal dependencies and narrative complexity, remains an open challenge.
@@
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (bai2025qwen25vltechnicalreport), Qwen2-VL (wang2024qwen2vlenhancingvisionlanguagemodels), and LongVU (wu2024longvideobench), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (liang2024tvrrankingdatasetrankedvideo; lei2021qvhighlightsdetectingmomentshighlights), action classification (caba2015activitynet; kinetics-400; soomro2012ucf101dataset101human), video reasoning (xie2024funqasurprisingvideocomprehension; xiao2021nextqanextphasequestionansweringexplaining), and video QA (xu2017video; lei2019tvqa; jang2017tgifqa). Among longer video benchmarks, TVQA (lei2019tvqa) is an early example with over 152.5 K QA pairs across 21.8 K clips (1.27 min avg.), totaling 461.2 hours. More recent datasets like MovieChat-1K(song2023moviechat) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(ates2020craft) and GazeVQA (ilaslan2023gazevqa) extend to causal and embodied reasoning. Other long-form efforts include MLVU (zhou2024mlvucomprehensivebenchmarkmultitask), MoVQA (zhang2023movqa), and Video-MME (videomme), though their average durations peak at 17 minutes. LVBench(wang2024lvbench) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(mangalam2023egoschema) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count in terms of long video understanding (87.7K).